### v1

@@ Abstract
Generative models, such as diffusion models (DMs), variational autoencoders (VAEs), and generative adversarial networks (GANs), produce images with a level of authenticity that makes them nearly indistinguishable from real photos and artwork. While this capability is beneficial for many industries, the difficulty of identifying synthetic images leaves online media platforms vulnerable to impersonation and misinformation attempts. To support the development of defensive methods, we introduce ImagiNet, a high-resolution and balanced dataset for synthetic image detection, designed to mitigate potential biases in existing resources. It contains 200K examples, spanning four content categories: photos, paintings, faces, and uncategorized. Synthetic images are produced with open-source and proprietary generators, whereas real counterparts of the same content type are collected from public datasets. The structure of ImagiNet allows for a two-track evaluation system: i) classification as real or synthetic and ii) identification of the generative model. To establish a baseline, we train a ResNet-50 model using a self-supervised contrastive objective (SelfCon) for each track. The model demonstrates state-of-the-art performance and high inference speed across established benchmarks, achieving an AUC of up to 0.99 and balanced accuracy ranging from 86% to 95%, even under social network conditions that involve compression and resizing. Our data and code are available at https://github.com/delyan-boychev/imaginet.
Keywords:
Synthetic Image Detection Contrastive Learning
1 Introduction
@@
State-of-the-art generative models are rapidly improving their ability to produce nearly identical images to authentic photos and artwork. Diffusion models (DMs) [19, 41], variational auto-encoders (VAEs) [17], and generative adversarial networks (GANs) [15] are being utilized in various ways to achieve data augmentation, text-to-image and image-to-image generation, inpainting and outpainting. They facilitate the production of visuals and spatial effects for downstream use in the entertainment, gaming, and marketing industries. On the other hand, these models can be misused by malicious actors. For instance, deceiving content and fake news [32] can be easily generated. Also, the flood of synthetic images and videos overwhelms moderation in content delivery platforms and social networks. Such issues are typically challenging and time-consuming to be solved manually. Thus, there is an increasing demand for improved synthetic image recognition models. Previous research [52, 16, 12] utilizes standard classifier training approaches to produce models with adequate performance on the task. They [16, 12] rely on a dataset with a single source of synthetic examples and a no-downsampling architecture to improve the performance of the detectors on images from unseen generators. These approaches do not completely solve the highly complex task of synthetic content detection due to overfitting, unreliability to detect novel generators, or biases toward various content types. Thus, important aspects of synthetic detection that have yet to be fully explored are the curation of a training dataset with a more diverse selection of content types and the utilization of training techniques better suited for generalizable detectors.
Figure 1: Examples from each category of the ImagiNet dataset.
We propose a new benchmark and balanced training set for synthetic image detection called ImagiNet (Figure 1). It includes images from state-of-the-art open-source and proprietary generators. Our main goal is to solve the problem with generalizability by training on diverse data. The images are created by either GAN [15], DM [40], or a proprietary generator – Midjourney [20] or DALL·E [6]. Our benchmark has two main testing tracks – synthetic image detection and model identification. The testing is conducted under perturbations such as JPEG compression and resizing, simulating social network conditions similarly to previous works [12]. For more consistent results, all the images are high-resolution, like the ones posted on social networks.
As discussed in other works, contrastive learning methods for pre-training detectors can achieve state-of-the-art performance on synthetic content detection. Wu et al. [52] utilize the textual embedder CLIP [38] to contrast image representations and class descriptions (e.g. "real photo", "synthetic painting") embeddings. However, it makes the training procedure computationally expensive and rather slow. Instead, we provide a more efficient contrastive baseline model for our dataset. The training paradigm we examine is called Self-Contrastive Learning (SelfCon) [4]. It is a modification of the Supervised Contrastive Learning method [25], which requires a multi-viewed (twice augmented) batch to improve the performance by allowing contrast of a single image within a batch. SelfCon provides a multi-exit architecture consisting of a standard network and a sub-network that can be attached to the intermediate layers. Its function is to produce variations of the embeddings in the latent space, acting as data augmentation. The training procedure has two stages: backbone pre-training and classifier training. The approach is suited to the structure of our dataset and allows for solving multiple tasks during training - real or synthetic classification and model identification.
2 Related Work
2.1 Previous Synthetic Datasets
Previous training and evaluation datasets (Table 1) present a large number of images for training and testing. Many of them include primarily images generated by GANs. However, contemporary generation approaches, such as DMs, often provide advantages in resolution and quality [13] compared to GANs. Thus, training on datasets limited to low-resolution images or primarily focusing on a specific set of generators [12, 16], often excluding proprietary models [52], can lead to limitations in the performance when detecting images from other unseen generators. Another significant problem is the lack of multiple content types during the training procedure, which makes the models potentially prone to overfitting. All these structural problems contribute to the introduction of biases [50].
Rahman et al. [39] propose a great alternative for benchmarking synthetic image detectors because of the high number of generators included and the diverse content types. However, neither a training set nor baseline weights are provided for comparison and benchmarking, limiting the dataset only for evaluation. Additionally, the dataset’s low-resolution images and the lack of balance may further constrain its application in the area.
Table 1: Feature comparison of previous synthetic datasets.. "-" signifies that data is not available. The maximum resolutions are mentioned below.
Training/Benchmark Corvi2022 [12] Wu2023 [52] ArtiFact [39] Ours
@@ 2.1 Previous Synthetic Datasets
2.2 Contrastive Learning
Contrastive Learning [11] is a training paradigm that produces generalizable models and is able to solve multiple representation learning tasks with one training objective – characteristics we find useful in synthetic image detection [53]. Optimizing contrastive objective clusters the representation vectors of training examples in the latent space.
The objective can be self-supervised, semi-supervised, or supervised and is inspired by N-pair losses [47]. In the self-supervised approach [9], an augmented copy of an input produces a representation (positive example) to be pulled closer to the anchor (original input) representation. The remaining images and their augmentations (negative examples) are pushed away in the latent space. In the semi-supervised and supervised approaches, images labelled as a given class and their augmented versions are included in the positive examples.
The contrastive learning involves pre-training a backbone encoder $F(.)$, e.g. ResNet-50 [18], which consists of feature-extraction convolutional layers. The encoder is attached to a Multilayer Perceptron [34] with Rectified Linear Unit (ReLU) [26] activation. This part of the architecture is referred to as a projection head $H(.)$, which outputs the normalized embeddings in $l_{2}$ space. It has been shown that without this component, the performance of the network degrades significantly [9].
For each batch, we sample $N$ examples and augment them twice to get $2N$ inputs $A\equiv\{1,...,2N\}$. We denote the first image of each pair as $x_{i}$ – anchor – and the second one as $x_{j(i)}$ – positive example. The anchor is contrasted with the other $2(N-1)$ data points. The output of the network for a given input image is $\omega(x)=H(F(x))$. Therefore the contrastive loss for all positive pairs is as follows:
$\mathcal{L}_{SelfSup}=-\sum\limits_{i\in A}\log\frac{\mathrm{exp}(\omega(x_{i})\cdot\omega(x_{j(i)})/\tau)}{\sum\limits_{l\in Q(i)}\mathrm{exp}(\omega(x_{i})\cdot\omega(x_{l})/\tau)}$ (1)
where $Q(i)\equiv A\backslash\{i\}$ (similarity between $z_{i}$ and $z_{i}$ is redundant), and $\tau$ is a temperature parameter. The dot product yields the cosine similarity between the vectors, as they are normalized in $l_{2}$ space.
2.2.1 Supervised Contrastive Learning
However, the self-supervised loss is impractical when we have labels $\hat{y}$ for all images in the batch. The Supervised Contrastive Loss (SupCon) [25] allows us to pull together all positive examples $P(i)\equiv\{p\in Q(i):\hat{y}_{p}=\hat{y}_{i}\}$ for a given anchor $x_{i}$ and push the rest of the negatives. With these terms added, we achieve the Supervised Contrastive loss:
$\mathcal{L}_{SupCon}=\sum\limits_{i\in A}\frac{-1}{\lvert P(i)\rvert}\sum\limits_{p\in P(i)}\log\frac{\mathrm{exp}(\omega(x_{i})\cdot\omega(x_{p})/\tau)}{\sum\limits_{l\in Q(i)}\mathrm{exp}(\omega(x_{i})\cdot\omega(x_{l})/\tau)},$ (2)
where we compute the loss for each example as an anchor and the average for each positive example of the given anchor.
2.2.2 Self-Contrastive Supervised Learning
Self-Contrastive Learning [4] offers a twice faster alternative to SupCon. A sub-network is attached to the backbone. Its main responsibility is to produce an alternative view of the images in the latent space instead of additional augmented samples to design the SelfCon loss with a single-viewed batch (augmented once). The sub-network could be a fully connected layer or another architecture with the same function as the backbone. The sub-net $H_{sub}(.)$ is attached to the backbone and projects the latent representations $F_{m}(.)$ obtained after the $m$-th ResNet block. The network has two output mapping functions $\Omega\equiv\{H_{sub}(F_{m}(.)),H(F(.))\}$ for a given input $x_{i}$. By introducing this set, we achieve each vector pair (from the main network and the sub-network) for a given input but with $A\equiv\{1,...,N\}$ because this method requires only a single augmentation:
$\mathcal{L}_{SelfCon}=\sum\limits_{\begin{subarray}{c}i\in A,\\ \omega\in\Omega\end{subarray}}\frac{-1}{\lvert P(i)\rvert\lvert\Omega\rvert}\sum\limits_{\begin{subarray}{c}p\in P(i),\\ \omega^{\prime}\in\Omega\end{subarray}}\log\frac{\mathrm{exp}(\omega(x_{i})\cdot\omega^{\prime}(x_{p})/\tau)}{\sum\limits_{l\in Q(i)}\mathrm{exp}(\omega(x_{i})\cdot\omega^{\prime}(x_{l})/\tau)}$ (3)
After the pre-training, a calibration stage is performed. A classifier is fitted to the representations generated by the pre-trained backbone encoder. The encoder’s weights remain frozen during this stage, allowing the classifier to adapt to the established latent distribution.
3 Methods
3.1 Dataset Construction
@@
ImageNet [43] 7.5K StyleGAN-XL [44] 7.5K
LSUN [54] 7.5K ProGAN* [21] 7.5K
COCO [27] 15K SD v2.1 [40]/SDXL v1.0 [37] 15K
@@
StyleGAN3 [23] 11.25K
SD v2.1 [40]/SDXL v1.0 [37] 5.625K
WikiArt [48] Danbooru [3] 11.25K 11.25K Animagine XL [49] 5.625K
@@
StyleGAN-XL [44] 11.25K
FFHQ [24] 22.5K SD v2.1 [40]/SDXL v1.0 [37] 11.25K
@@
Midjourney* [20] 12.5K
Photozilla [46] 25K DALL·E 3* [6] 12.5K
@@
The ImagiNet dataset consists of images from various open-source and proprietary image generators to encompass the distinct "fingerprints" they impart. Dataset’s name combines imagine, an often-used instruction in text-to-image generation, with the status quo naming of visual databases used in computer vision. Imagination also emphasizes that we focus on the content type diversity of the images generated through creative prompts.
@@
Dataset Structure (Table 2) – The dataset structure is designed to represent real-world scenarios where images of different content types might be used. ImagiNet examples are split into two main categories – real and synthetic images. To mitigate content-related biases, the dataset is divided into four subcategories—photos, paintings, faces, and uncategorized. Such images are commonly found on the World Wide Web and are the main subject of generative applications. We provide a balanced amount of synthetically generated images and real counterparts in each subcategory. The source datasets and generator models are given in Table 2. Images from models marked with * are sourced as follows: ProGAN from Wang et al. [51], Midjourney from Pan et al. [35], DALL·E 3 from LAION [2]; in addition we generated 800 DALL·E 3 images to reach our desired dataset size by prompting it through the OpenAI API. The unmarked synthetic groups are generated automatically using the respective pre-trained models. Samples from all GAN models are labelled with the GAN category, and samples from the Stable Diffusion variants are labelled as diffusion for the model detection task. Proprietary models are labelled as standalone categories. A detailed analysis of the presence of multiple content types and synthetic generators is available in the supplementary materials.
@@
Real Images Sampling – The real images are randomly sampled from each real counterpart dataset placed in Table 2. The images in our test set are sampled from the validation and testing parts of these sets.
@@
Image Generation Procedure – To generate images with GANs, we sample random latent code (it is selected according to model requirements) for a given seed and feed the generator with it unconditionally. For DMs and private generators (SD v2.1, SDXL v1.0, Animagine XL, DALL·E 3), however, textual guidance is needed, thus we first search manually for appropriate negative prompts and positive suffixes to increase the quality of the produced images. The construction of each prompt is in descriptive form. For photos, we utilize the captions from COCO [10] to prompt the generators and achieve images with sufficient quality. For paintings, instead of using a pre-generated set of captions for prompting, we create lists of styles, techniques, and subjects with GPT-3.5 Turbo [7]. After that, we fit these characteristics of the paintings in a descriptive sentence shown in Figure 2(a), which guides the model to generate varied images. The gaps are filled respectively with an item from the given list, and in the end, a positive suffix is added. The procedure for face generation is similar – Figure 2(b) presents the structure of the prompt. All the lists for filling in the guiding instructions, as well as the positive suffixes and negative prompts, are available in our supplementary material. The last model AnimagineXL, a fine-tuned SDXL [37] variant for art generation, uses only tags from the Danbooru dataset [3] for prompting.
@@
Dataset Splits – From the whole set, we sample 80% of the images from each category and subcategory with an equal number of images from the different generators. The number of images in the training set is 160K, respectively 40K are left for testing. Because of the contrastive approach that we apply, we need a balanced calibration set sampled from the training set (an equal number of images for each model). It consists of 80K examples in total.
Labellings and Evaluation Tracks – All the images of the dataset are labelled. They have four labels – source (real or synthetic), content type, generator group (GAN, SD, etc.), and specific generator. In our benchmark, we have two tracks – synthetic image detection and model identification. On the test set are applied perturbations to simulate social network conditions [12]. First, we do a large square crop (ranging from 256 to the smaller dimension of the image) of the image and, after that, resize it to $256\times 256$. After that, we compress 75% of the images with JPEG or WebP compression.
@@
3.2 Training Procedure
@@
To train our model, we first initialize the ResNet-50 model with pre-trained ImageNet weights and modify its early layers to avoid downsampling, following Gragnaniello et al. [16]. In the first stage, we train the backbone with a compound contrastive objective $\mathcal{L}_{SelfCon}$. In our case, the mapping functions $H(.)$ and $H_{sub}(.)$ output representations in $\mathbb{R}^{128}$. This involves accumulating $\mathcal{L}_{SelfCon}$ applied on two labellings - synthetic detection and model identification labels, with each loss assigned equal weight. When optimizing the model detection objective, real images in the batch are ignored. To address the increased memory demands of removing downsampling in early ResNet-50 layers and the large batch size requirements of SelfCon, we adopt gradient caching [14], a technique initially designed for language model contrastive losses. We modify [^1] Our implementation of gradient caching for contrastive objectives is available at https://github.com/delyan-boychev/grad-cache-con-learning it for use with SelfCon [4], SupCon [25], and SimCLR [9]. This approach calculates the loss on the entire batch but accumulates gradients in smaller chunks, allowing for large batch sizes and efficient training on memory-constrained GPUs.
@@
The second stage involves calibrating the model. We detach the sub-network and projection heads, replacing the output projection head with a multilayer perceptron classifier. This classifier is then trained using cross-entropy loss on a balanced dataset to perform both origin and model detection. We update the batch normalization statistics within the backbone’s residual blocks, following Schneider et al. [45], to enhance robustness against perturbations not encountered during pre-training.
4 Experiments and Results
We provide a single model following the procedure explained in Section 3.2. During the first stage, the backbone is optimized with SGD [42] for 400 epochs with batch size $N=200$ on the ImagiNet training set. The initial learning rate of $0.005$ is warmed up linearly [31] for 10 epochs and is cosine annealed [29] afterwards. The second stage continues for 5 epochs with AdamW optimizer [29] and constant learning rate $0.0001$, weight decay $0.001$, $\beta_{1}=0.9$ and $\beta_{2}=0.99$. After the pre-training procedure, we visualize the model representations of the images in the test set by applying Autoencoder dimensionality reduction [33] (discussed in supplementary materials). We obtain the plots in Figure 3 with our model’s representations computed on center-cropped images from our test set. The plots show the ability of our model to distinguish between real and synthetic images as well as to cluster each generator’s images.
Additionally, we train a model on the ImagiNet training set utilizing the Corvi et al. [12] approach and fine-tune the classifier to have a fair comparison with our novel method. This calibration method did not result in a statistically significant improvement in either accuracy or AUC on average across different benchmarks  [^2] The fine-tuned classifier is not included in the comparison, as its performance is statistically indistinguishable from the presented model..
(a) Synthetic Detection
(b) Model Detection
Figure 3: Dimensionality reduction plots of the representations of the model computed on a part of the ImagiNet test set.
4.1 Synthethic Detection Track
Testing Settings – In our evaluation report, we showcase our model and four other approaches - Grag2021 [16] trained on StyleGAN images, Corvi2022 [12] (where logits of two classifiers are fused), Wu2023 [52], and Corvi2022 [12] trained on ImagiNet training set (a single model). The raw probabilities are computed following the requirements of each model. Wu2023 detector leverages cosine similarity for predictions. To accommodate this approach, an anchor vector is required, which is obtained from real images by averaging their latent representations at the final layer of the network. For model-to-model comparability, we derive a distinct anchor vector for each test set, estimated using 100 real images drawn from the corresponding benchmark. Furthermore, the predicted probabilities of the detectors are calibrated with the Platt scaling method [36] by sampling 2 real and 2 synthetic images from each generator in the given test set. For the ImagiNet benchmark, we sample 10 real and 10 synthetic to mitigate potential biases since the GAN and SD categories contain different approaches to generative modeling. We evaluated the performance of all approaches against two established benchmarks: The practical test set (Table 3) and Corvi’s test set (Table 4). We also tested the detectors on our novel benchmark – ImagiNet (Table 5). The practical test is designed to assess detection performance under ideal conditions. It utilizes $448\times 448$ crops extracted from the original image. Corvi’s and ImagiNet benchmarking are conducted with a random large crop, followed by a resize to $256\times 256$ and Jpeg or WebP compression on the images to simulate social network scenarios.
[…]

### v2

@@ Abstract
Recent generative models produce images with a level of authenticity that makes them nearly indistinguishable from real photos and artwork. Potential harmful use cases of these models, necessitate the creation of robust synthetic image detectors. However, current datasets in the field contain generated images with questionable quality or have examples from one predominant content type which leads to poor generalizability of the underlying detectors. We find that the curation of a balanced amount of high-resolution generated images across various content types is crucial for the generalizability of detectors, and introduce ImagiNet, a dataset of 200K examples, spanning four categories: photos, paintings, faces, and miscellaneous. Synthetic images in ImagiNet are produced with both open-source and proprietary generators, whereas real counterparts for each content type are collected from public datasets. The structure of ImagiNet allows for a two-track evaluation system: i) classification as real or synthetic and ii) identification of the generative model. To establish a strong baseline, we train a ResNet-50 model using a self-supervised contrastive objective (SelfCon) for each track which achieves evaluation AUC of up to 0.99 and balanced accuracy ranging from 86% to 95%, even under conditions that involve compression and resizing. The provided model is generalizable enough to achieve zero-shot state-of-the-art performance on previous synthetic detection benchmarks. We provide ablations to demonstrate the importance of content types and publish code and data.
Introduction
@@
State-of-the-art generative models are rapidly improving their ability to produce nearly identical images to authentic photos and artwork. Diffusion models (DMs) (Ho, Jain, and Abbeel 2020; Rombach et al. 2022a), variational auto-encoders (VAEs) (Harvey, Naderiparizi, and Wood 2022), and generative adversarial networks (GANs) (Goodfellow et al. 2014) are being utilized in various ways to achieve data augmentation, text-to-image and image-to-image generation, inpainting and outpainting. They facilitate the production of visuals and spatial effects for downstream use in the entertainment, gaming, and marketing industries. On the other hand, these models can be misused by malicious actors (Masood et al. 2021). Thus, there is an increasing demand for improved synthetic image recognition models. Prior work (Wu, Zhou, and Zhang 2023; Gragnaniello et al. 2021; Corvi et al. 2022) employs standard classifiers but struggles with overfitting, bias, and poor generalization to novel generators, limiting effectiveness in synthetic content detection. One key area that has yet to be fully explored in synthetic detection is the creation of training datasets with a broader range of content types and generator sources.
Previous datasets (Table 1) primarily feature GAN-generated images and lack diversity in resolution, generator types, and content, leading to biases and overfitting issues (Corvi et al. 2022; Gragnaniello et al. 2021; Wu, Zhou, and Zhang 2023; Torralba and Efros 2011). Rahman et al. (2023) provide a diverse benchmark with multiple generators and content types, but the resized low-resolution images make it more suitable for benchmarking rather than training.
Train/Eval Corvi2022 Wu2023 ArtiFact Ours
@@
Table 1: Feature comparison of previous synthetic datasets. ‘-’ signifies that data is not available.
We propose a new benchmark and balanced training set for synthetic image detection called ImagiNet111 https://github.com/delyan-boychev/imaginet. It includes images from novel open-source and proprietary generators. Our main goal is to study ways to address the challenge of generalizability by training on diverse data. The images are created by either GAN (Goodfellow et al. 2014), DM (Rombach et al. 2022b), or a proprietary generator – Midjourney (Holz 2023) or DALL·E (Betker et al. 2023). Our benchmark includes two main testing tracks: synthetic image detection and model identification. Testing is performed under perturbations like JPEG compression and resizing, simulating social network conditions as in previous works (Corvi et al. 2022). All images are high-resolution, similar to those on social networks, for more consistent results.
Dataset Construction
The ImagiNet dataset consists of images from various open-source and proprietary image generators to encompass the distinct “fingerprints” they impart.
@@
ImageNet 7.5K StyleGAN-XL 7.5K
LSUN 7.5K ProGAN* 7.5K
COCO 15K SD v2.1/SDXL v1.0 15K
@@
StyleGAN3 11.25K
SD v2.1/SDXL v1.0 5.625K
WikiArt Danbooru 11.25K 11.25K Animagine XL 5.625K
@@
StyleGAN-XL 11.25K
FFHQ 22.5K SD v2.1/SDXL v1.0 11.25K
@@
Midjourney* 12.5K
Photozilla 25K DALL·E 3* 12.5K
@@
Dataset Structure (Table 2) – The dataset structure is designed to represent real-world scenarios where images of different content types might be used. ImagiNet examples are split into two main categories – real and synthetic images. To mitigate content-related biases, the dataset is divided into four subcategories – photos, paintings, faces, and miscellaneous. Such images are commonly found on the World Wide Web and are the main subject of generative applications. We provide a balanced amount of synthetically generated images and real counterparts in each subcategory. The source datasets and generator models are given in Table 2. Images from models marked with * are sourced as follows: ProGAN from Wang et al. (2020), Midjourney from Pan et al. (2023), DALL·E 3 from LAION (LAION 2023); in addition we generated 800 DALL·E 3 images to reach our desired dataset size. Synthetic groups are generated with pre-trained models: GAN images are labeled as GAN, Stable Diffusion as SD, and proprietary models as standalone.
@@
Real Images Sampling – The real images are randomly sampled from each real counterpart dataset described in Table 2 (Russakovsky et al. 2015; Yu et al. 2016; Lin et al. 2015; Tan et al. 2019; Anonymous, community, and Branwen 2022; Karras, Laine, and Aila 2019; Singhal et al. 2021). The images in our test set are sampled from the validation and testing splits of these sets.
@@
Image Generation Procedure – To generate images with GANs (StyleGAN-XL (Sauer, Schwarz, and Geiger 2022), StyleGAN3 (Karras et al. 2021)), we sample random latent code (it is selected according to model requirements) for a given seed and feed the generator with it unconditionally. For DMs and private generators (SD v2.1 (Rombach et al. 2022b), SDXL v1.0 (Podell et al. 2023), Animagine XL (Taqwa 2024), DALL·E 3 (Betker et al. 2023)), however, textual guidance is needed, thus we first search manually for appropriate negative prompts and positive suffixes to increase the quality of the produced images. The construction of each prompt is in descriptive form. For photos, we utilize the captions from COCO (Chen et al. 2015) to prompt the generators and achieve images with sufficient quality. For paintings, instead of using a pre-generated set of captions for prompting, we create lists of styles, techniques, and subjects with GPT-3.5 Turbo (Brown et al. 2020). After that, we fit these characteristics of the paintings in a descriptive sentence shown in Figure 1(a), which guides the model to generate varied images. The gaps are filled respectively with an item from the given list, and in the end, a positive suffix is added. The procedure for face generation is similar – Figure 1(b) presents the structure of the prompt. All the lists for filling in the guiding instructions, as well as the positive suffixes and negative prompts. The last model AnimagineXL, a fine-tuned SDXL (Podell et al. 2023) variant for art generation, uses only tags from the Danbooru dataset (Anonymous, community, and Branwen 2022) for prompting.
@@
Dataset Splits – From the whole set, we sample 80% of the images from each category and subcategory with an equal number of images from the different generators. The number of images in the training set is 160K, respectively 40K are left for testing. We aim to provide a balanced (an equal number of images for each model) calibration set sampled from the training set. It consists of 80K examples in total.
Labelling and Evaluation Tracks – All the images of the dataset are labelled. They have four labels – source (real or synthetic), content type, generator group (e.g., GAN), and specific generator (e.g., ProGAN). In our benchmark, we have two tracks – synthetic image detection and model identification. Perturbations are applied on the test set to simulate social network conditions (Corvi et al. 2022). First, we do a large square crop (ranging from 256 to the smaller dimension of the image) of the image and, after that, resize it to $256\times 256$. After that, we compress 75% of the images with JPEG or WebP compression.
@@
Baseline Training
To train our baseline, we initialize a ResNet-50 model with pre-trained ImageNet weights and modify its early layers to avoid downsampling, following Gragnaniello et al. (2021).
In the first stage of training, we train a backbone with a contrastive objective $\mathcal{L}_{SC}$, as proposed by Bae et al. (2022):
$\mathcal{L}_{SC}$ $=\sum\limits_{\begin{subarray}{c}i\in A\\ \omega\in\Omega\end{subarray}}\frac{-1}{\lvert P(i)\rvert\lvert\Omega\rvert}$
$\quad\sum\limits_{\begin{subarray}{c}p\in P(i)\\ \omega^{\prime}\in\Omega\end{subarray}}\log\frac{\mathrm{exp}(\omega(x_{i})\cdot\omega^{\prime}(x_{p})/\tau)}{\sum\limits_{l\in Q(i)}\mathrm{exp}(\omega(x_{i})\cdot\omega^{\prime}(x_{l})/\tau)}$ (1)
where $A\equiv\{1,...,N\}$ is a set of indices for all batch examples, $Q(i)\equiv A\backslash\{i\}$ (similarity between $z_{i}$ and $z_{i}$ is redundant), and $P(i)\equiv\{p\in Q(i):\hat{y}_{p}=\hat{y}_{i}\}$ is the set of positive examples for a given example $i$.
@@
A sub-network is attached to the backbone. Its main responsibility is to produce an alternative view of the images in the latent space instead of additional augmented samples to design the SelfCon loss with a single-viewed (augmented once) batch. The sub-network could be a fully connected layer or another architecture with the same function as the backbone. The sub-net $H_{sub}(.)$ is attached to the backbone and projects the latent representations $F_{m}(.)$ obtained after the $m$-th ResNet block. The network has two output mapping functions $\Omega\equiv\{H_{sub}(F_{m}(.)),H(F(.))\}$ for a given input $x_{i}$. In our case, the mapping functions $H(.)$ and $H_{sub}(.)$ output representations in $\mathbb{R}^{128}$. This involves accumulating $\mathcal{L}_{SC}$ applied on two labellings - synthetic detection and model identification labels, with each loss assigned equal weight. When optimizing the model detection objective, real images in the batch are ignored. To address the increased memory demands of removing downsampling in early ResNet-50 layers and the large batch size requirements of SelfCon, we adopt gradient caching (Gao et al. 2021), a technique initially designed for language model contrastive losses. We modify it for use with SelfCon (Bae et al. 2022), SupCon (Khosla et al. 2021), and SimCLR (Chen et al. 2020). This approach calculates the loss on the entire batch but accumulates gradients in smaller chunks, allowing for large batch sizes and efficient training on memory-constrained GPUs.
@@
The second stage involves calibrating the model. We detach the sub-network and projection heads, replacing the output projection head with a multilayer perceptron classifier. This classifier is then trained using cross-entropy loss on a balanced dataset to perform both origin and model detection. We update the batch normalization statistics within the backbone’s residual blocks, following Schneider et al. (2020), to enhance robustness against perturbations not encountered during pre-training.
Experiments and Results
Figure 2: Dimensionality reduction vizualization of the backbone representations for a subset of ImagiNet.